Question 1 (1 point) Given the frequency table below for a list of data, find the mean of the data. Enter the number only In your answer. For example, if you found the mean was 4, you would enter 4. Value Frequency 9 3 2 3 2 6 2 10 12 13 14 Provide your answer below:
Question 2 (1 point) Find the mode of the following list of numbers. 10,10,7,18,11,14,7,11,14,11,18 Give your answer as a number only. For example, if you found the mode was 9. you would enter 9. Provide your answer below: e® 1-11 Paragraph Previous Page Next Page Page 2 co
media%2F314%2F314636ec-d269-4f00-a60e-82
For a certain data set, the minimum is 4, the median is 16, the third quartile is 19, the interquartile range is 4, and the maximum is 20. Which of these shows the correct box-and-whisker plot for the data? 0 2 4 6 8 10 12 14 16 18 20
Answer #1

Solution:

Question 1:

The mean is given below:

Frequency Value x 32 27 20 4 9 2 10 12 13 14 2 6 2 24 78 28 yfX = 242 Σί-22 Therefore, the mean is: yfc 242 Mean 11

Question 2:

The mode is the most frequently occurring observation in the data set. We clearly see 11 occurs maximum 3 times in the given data set. Therefore, the mode = 11.

Question 3:

What is the probability of a family owning a cat or a dog?

Answer:

Plcat-or-dog) = P(cat) + P(dog)-P(cat-and-dog)

  - 0.1 + 0.2 - 0.08 - 0.22

Therefore, the option 0.22 is correct

For a certain data set, the minimum is 4, the median is 16, the third quartile is 19, the interquartile range is 4, and the m

First quartile = Third quartile - IQR = 19 - 4 = 15